OpenStudy (anonymous):

each question of a 5 multiple exam has 4 possible answers. suppose a student picks an answer at ramdon for each question.find the prob the student slects the correct answer on only the 2nd and 3rd questions

OpenStudy (dumbcow):

= 3/4 * 1/4 *1/4 * 3/4*3/4 =(1/4)^2 *(3/4)^3 =(1/16)*(27/64) =27/1024

OpenStudy (anonymous):

so 1024?

OpenStudy (dumbcow):

no probabilities are between 0 and 1 p = 27/1024

OpenStudy (anonymous):

that would be the answer of u have to deveide

OpenStudy (dumbcow):

if you divide it its .026367188

OpenStudy (anonymous):

would u have to devide?

OpenStudy (dumbcow):

the answer in fraction form is 27/1024 the answer in decimal form is 0.026367188

OpenStudy (anonymous):

which one would i use because it says problility

OpenStudy (dumbcow):

it doesn't matter since it's the same answer but typically for probability questions they want the answer in the fraction form

OpenStudy (anonymous):

HELP ME I JUST POSTED A NEW ONE